OpenStudy (anonymous):

given the equation 5x-3y=15, how would you graph the line USING the x and y int.?

OpenStudy (anonymous):

would u just simply put a point on -3 and go up 5 and run 1?

OpenStudy (agent0smith):

5x-3y=15 To find y-int. plug in x=0, solve for y To find x-int: plug in y=0, solve for x

OpenStudy (agent0smith):

Once you do that, it's easy to graph.

OpenStudy (anonymous):

thanks!!!

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

it just said to graph 5x-3y=15 by using the x and y int. so i just write 5(3)-3(5)=15?

OpenStudy (anonymous):

.-.

OpenStudy (agent0smith):

No. Do exactly as i said above FIRST.

OpenStudy (anonymous):

OH! so x=3 and y=5?

OpenStudy (agent0smith):

x-int is correct, y isn't.

OpenStudy (anonymous):

isnt y correct because its : 5(0)-3y=15

OpenStudy (anonymous):

ohhhhh its -5

OpenStudy (agent0smith):

now you have your x and y intercepts. Use them to plot the graph.

OpenStudy (anonymous):